OpenStudy (anonymous):

Polar coordinates. z=xy and 0

OpenStudy (anonymous):

It looks as though one of your variables is a bit goofed. My presumption would be that it should be theta that runs from 0 to pi/2. In that case, x = r cos theta and y = r sin theta.

OpenStudy (anonymous):

yes yes... sorry typo

OpenStudy (anonymous):

0<r<1

OpenStudy (anonymous):

you went trough a integration by parts right?

OpenStudy (anonymous):

\[xy= r \sin(\theta ) r \cos(\theta) = \frac{1}{2} r^2 \sin(2\theta) \]

OpenStudy (anonymous):

The jacobian of the plar substitution is r

OpenStudy (anonymous):

\[I = \frac{1}{2}\int\limits_{0}^{1} r^3 dr \int\limits_{0}^{\frac{\pi}{2}} \sin (2\theta ) d \theta \]

OpenStudy (anonymous):

yesss thats it :P thanks!

OpenStudy (anonymous):

ugh im new to this , wheres the good answer button?

OpenStudy (anonymous):

\[= \frac{1}{2} \times \frac{1}{4} \times 1 = \frac{1}{8}\]

OpenStudy (anonymous):

just to the right of the username

OpenStudy (anonymous):

right to ur username theres "group champion" and a gold medal followed by a zero, theres no actual button to click

OpenStudy (anonymous):

huh, strange , ahh dw bout it

OpenStudy (anonymous):

there is mean to be a blue button right next to the medal thing

OpenStudy (anonymous):

nope... im using safari i dont know if thats something to it ...

OpenStudy (anonymous):

yup, safari bugged , there you go. thank you again ;)